# Imprints of Supermassive Black Hole Evolution on the Spectral and Spatial Anisotropy of Nano-Hertz Stochastic Gravitational-Wave Background

### v1

@@ 7 Results: Isotropic and Anisotropic SGWB from simulations
For the case where $z_{0}$ = 1, most of the sources are sampled from the coalescence rate distribution that decreases with frequency. In Fig. 6, we show the number of sources emitting in each frequency bin ($\Delta f$) and each redshift bin ($\Delta z$ = 0.1), represented by a color bar, along with the theoretical curve at 3 different redshifts for scenarios of the coalescence rate case. It can be seen that the slope of the $\frac{dN_{f}}{df}\Delta f$ decreases with redshift for redshift-dependent scenarios (Fig. 6 and Fig. 6) whereas it remains stable in the redshift independent scenarios (Fig. 6 and Fig. 6) of the coalescence rate, as expected. For the redshift-dependent case, the number of sources at low redshift in high-frequency bins becomes so small that the expected number of sources in that frequency bin at that redshift is below 1. However, the Monte Carlo sampling only returns the discrete integer source. Therefore, the analytical curve lies below the simulated results at high frequencies. The effect of this is also evident in the $C_{\ell}(f)$ for two cases: when $\delta_{f}=0$ and $\mu=0$, and when $\delta_{f}=1$, $\mu=1$, and $z_{0}=1$, where there is a sudden jump at $f=2.4\times 10^{-8}$. Similar to the $\nu=0.5$ curve in Fig. 8, the expectation value of the number of sources emitting in high-frequency bins from low redshift bins is less than one, but the sampled number of sources is mostly either zero or one. This sporadic sampling at high-frequency bins results in a sharp increase in the anisotropy at high frequencies.
The effect of this is also seen in the $C_{\ell}(f)$ for two cases: when $\delta_{f}=0$ and $\mu=0$, and when $\delta_{f}=1$, $\mu=1$ and $z_{0}=1$, where there is a sudden jump at $f=2.4\times 10^{-8}$. Just like in the $\nu=0.5$ curve in Fig. 8, the expectation value of the number of sources in the low redshift bins, emitting at high frequencies bins, is less than one, but the sampled number of sources can be mostly either zero or one. This sporadic nature of the sampling at high-frequency bins results in a sharp increase in the anisotropy at high frequencies.
@@
The fitted curves for $C_{\ell}(f)$ show a significantly larger power law index ($\lambda$) for the redshift-dependent case compared to the redshift-independent case for the same $\delta_{f}$ value. This difference arises because the number of sources in the redshift-dependent case decreases as the frequency increases. Similarly, the power law index is larger for $\delta_{f}=0$ than that of $\delta_{f}=1$ for the redshift independent case because there are fewer sources at high frequencies for the latter case than for the former.
@@
The redshift evolution of the coalescence rate can help us understand the merger history of the galaxies. A consistent coalescence rate across redshifts suggests a stable merger rate of galaxies over time. Conversely, if relatively more binaries are found at lower frequencies at some redshifts, it could indicate a higher merger rate during those periods. In the case where the coalescence rate at low frequency (compared to high frequency) increases with redshift, signifying a larger galaxy merger rate at high redshift, the anisotropy of such signal will have a larger slope as compared to the case where the coalescence rate is stable across redshifts.
@@
Variation of $C_{l}$ with spatial frequencies ($\ell$) and spectral frequencies($f$): In Fig. 11 we depict the spectral and spatial structure of the anisotropies in the SGWB for a few different models. The figure shows the mode of $C_{\ell}(f)$ over 1000 realizations as a function of the multipole moment ($\ell$) and frequency ($f$) in color. The results for different variations of the model parameters for each of these cases at a fixed spatial frequency $\ell=15$ are discussed previously. Here we mainly focus on the additional gain by measuring the signal in $\ell$ as well. The $\ell=0$ mode signifies the isotropic component ($\Omega_{gw}(f)$) of the SGWB signal. The presence of non-zero higher-order moments implies additional information in the $\ell-f$ space beyond just the $\ell=0$ (isotropic) moment. The non-zero higher-order moments indicate the presence of anisotropy in the SGWB as a function of GW frequency. As evident from the color matrix plots in Fig. 11, the $\ell-f$ structure remains nearly invariant with respect to $\ell$ at all frequencies ($f$). This is expected as at these large angular scales, the properties of the host galaxies and the correlation with the n-Hz GW signal will be similar for a statistically isotropic and homogeneous Universe.
@@
In Fig. 11– Fig. 11, we display the $\ell-f$ plot of the $C_{\ell}(f)$ for four different cases. Consistent with the behavior at $\ell=15$, the $C_{\ell}(f)$s curves exhibit an increase from approximately one to twelve as the frequency increases from $f=2\times 10^{-9}$ Hz to $f=3\times 10^{-8}$ Hz in nearly a steady way at all values of $\ell$. The steady increase at all $\ell$ for every spectral frequency $f$ makes the theoretical signal strongly correlated and predictable. As the Universe is expected to be statistically isotropic and homogeneous at large scales, the SGWB fluctuations at large scale are nearly independent of spatial scale $\ell$ values and vary strongly with the spectral frequency $f$ of GWs, which captures the evolution of the SMBHBs. If one can probe angular scales corresponding to galaxies and smaller (higher $\ell$s), the spatial behavior is expected to show some variation. As the current PTA analysis can only explore large angular, we have not explored the signal at small angular scales in this paper. The small angular scale signal will be explored in a future work.
@@
The presence of anisotropic power spectrum with more power at high $f$ for every value of spatial frequency $\ell$ makes it possible to measure the amplitude $B(l)$ and the power-law index $\lambda$ (of a two-parameter model to capture the anisotropic power spectrum) of $C_{l}(f)$ by combining different values of $\ell$, which gives a $\sqrt{f_{\rm sky}(2\ell+1)}$ independent information per $\ell$ for a survey with sky-fraction $f_{\rm sky}$. The theoretical analysis presented in this paper demonstrates how the anisotropic signal of the SGWB in the $\ell-f$ space can capture the imprints of SMBHBs evolution in the Universe. We have shown the variation of $C_{l}(f)$ for different scenarios of astrophysical models and have theoretical predictions on the scaling of the anisotropic signal with $f$. In future work, we will explore the measurability of this signal and how it can be constrained from the SGWB observations.
@@
A summary of the key signatures of SMBHBs evolution on the SGWB spatial-spectral frequency are:
@@
  • Even in the absence of any redshift evolution of the SMBH mass and galaxy stellar mass relation, the anisotropy in SGWB captured in terms of the spatial-spectral power spectrum $C_{l}(f)$ can be large if SMBH masses and galaxy stellar mass relation is more steep implying that heavier SMBHs can be efficiently hosted in galaxies of lower stellar mass. The amplitude and shape of SGWB power spectrum $\Omega_{\rm gw}(f)$ can be comparable for both steep and shallow SMBH masses and galaxy stellar mass relation.
@@
  • If heavier SMBHs form efficiently at high redshift, then the amplitude of the anisotropy can increase in the high frequencies of n-Hz signal than the low frequency for a comparable SGWB power spectrum $\Omega_{\rm gw}(f)$. The spectral shape of $\Omega_{\rm gw}(f)$ and spatial-spectral anisotropy power spectra $C_{l}(f)$ bring complementary information on the redshift evolution of the SMBH and galaxy stellar mass relation.
@@
  • The coalescing rate of SMBHBs across cosmic time shows prominent features in both power spectrum $\Omega_{\rm gw}(f)$ and spatial-spectral anisotropy power spectrum $C_{l}(f)$. If the coalescence rate is less efficient in high frequency than in low frequency, and/or is less efficient at low redshift, then the amplitude of $C_{l}(f)$ increases. By measuring $\Omega_{\rm gw}(f)$ and $C_{l}(f)$ the frequency and redshift evolution of the coalescing rate can be inferred using the n-Hz signal.
@@ 8 Conclusion
However, to truly comprehend the SMBH population and its evolutionary trends, it is important to explore additional observational avenues and develop comprehensive theoretical frameworks that can account for the diverse range of processes involved in their formation and growth. Our study has highlighted the importance of studying the isotropic as well as the anisotropic SGWB density in the study of the SMBH signal which can shed light on the formation scenarios of SMBHBs. The study of SGWB anisotropy holds great promise as a valuable tool for advancing our understanding of some of the open questions regarding the formation and evolution of SMBHs and SMBHBs. The evolutionary history of the SMBHs is expected to have an imprint on the correlation between the SMBH mass and the properties of the host galaxies. To capture the physics from multi-scale, we adopted a multiscale adaptive technique to cross-match the physics involved at different scales from cosmological to the sub-pc scale where the coalescence of the binary occurs, to calculate the large-scale spatial and spectral anisotropy of the SGWB in n-Hz range. By utilizing large-scale simulations such as MICECAT and smaller-scale, high-resolution simulations such as ROMULUS25, along with an analytical model for the physics of the sub-pc SMBH environment, we demonstrate the impact of SMBHB evolution on the SGWB signal.
Our results have illustrated how the anisotropy of the SGWB from the population of the SMBHBs is affected by the SMBH mass-stellar mass relation. Similarly, we have also demonstrated that the coalescence rate and its evolution affect the overall shape of the spatial dependence of the angular correlation $C_{\ell}(f)$ of the SGWB. The coalescence rate of the SMBHBs at different separations of binaries can be used as a probe to the merger history of the galaxy. Furthermore, our investigation highlights the significant impact of environmental effects on the anisotropy of the SGWB. The presence of environmental effects, particularly at large separations of the binary, can alter the residence time of a binary at low frequencies. This effect leads to an increase in the anisotropy.
@@
Although with the present n-Hz data, the sky localization is poor, the future inclusion of the Square Kilometre Array (SKA) will significantly improve the sky localization error, as a large number of pulsars can be detectable with SKA (Lazio, 2013; Stappers et al., 2018). SKA is projected to find several hundred pulsars, which will enhance the PTA’s ability to detect and localize the GW signal. The sky localization of the PTAs scales as the square root of the number of pulsars (Sesana & Vecchio, 2010a, b). Therefore, with the addition of hundreds of pulsars, SKA is expected to enhance the localization of the GW signal to a few tens of square degrees (Sesana & Vecchio, 2010a, b).
The approach proposed in the work will help us put some valuable constraints on the population and evolution of the SMBHs. Using the multiscale adaptive technique employed in this work, we can efficiently calculate the spatial-spectral anisotropy of the SGWB signal for a wide range of astrophysical scenarios of SMBH formation and evolution, which can be tested with n-Hz observations. In the future, incorporating this framework into a Bayesian data analysis setup will enable the measurement of parameters (such as $\rho,\eta,\nu,\delta_{f},\kappa,\mu,z_{0}$) by using the spatial-spectral signal $C_{l}(f)$. Furthermore, cross-correlation of the SGWB signal with the tracers of the large-scale structures (such as galaxies, quasars, AGN) will bring additional insights into the formation of the SMBHs in the Universe. A work is currently under preparation to explore these aspects. Though this study strictly limits the power spectrum of the SGWB, signatures beyond the power spectrum, such as non-Gaussian signatures are also expected to be important due to the sporadic spatial distribution of the SMBHBs in the Universe. These will also bring more constraining power in measuring the SMBHB’s evolution parameters.
@@ Appendix A Fitting power law to the 15-year data release of NANOGrav
We fit the $\Omega_{gw}(f)$ of the NANOGrav 15-yr data set to a power law using PTArcade code (Mitridate et al., 2023).
$\Omega_{gw}(f)=A\times(f/f_{yr})^{\alpha},$ (19)
where A is the amplitude of the $\Omega_{gw}(f)$ at $f=f_{yr}$, $f_{yr}$ = 1/yr, and $\alpha$ is the power law index. Fig. 12 shows the corner plot of A and $\alpha$ fitted to the NANOGrav 15-yr data set. The uncertainties on parameters based on the 16th, 50th, and 84th percentiles are $\rm Log_{10}$= $-7.14_{-0.29}^{+0.28}$ and $\alpha$= $1.76_{-0.34}^{+0.36}$. We have normalized the mode over 1000 realizations of the $\Omega_{gw}(f)$ in each frequency bin by the 50 % confidence level of amplitude A and power law index $\alpha$ for cases ’SMBH mass-Stellar mass’ and ’SMBH mass-Stellar mass evolution’ mentioned in the Table. 1.
Figure 12: Corner plot showing the posterior distributions of $\rm Log_{10}(A)$ and $\alpha$ for the power-law model fitted to $\Omega_{\rm gw}$ estimates from NANOGrav 15-yr data set.

### v2

@@ 6 Results: Isotropic and Anisotropic SGWB from simulations
For the case where $z_{0}$ = 1, most of the sources are sampled from the coalescence rate distribution that decreases with frequency. In Fig. 5, we show the number of sources emitting in each frequency bin ($\Delta f$) and each redshift bin ($\Delta z$ = 0.1), along with the theoretical curve at 3 different redshifts for different scenarios of the coalescence rate case. It can be seen that the slope of the $\frac{dN_{f}}{df}\Delta f$ decreases with redshift for redshift-dependent scenarios (Fig. 5 and Fig. 5) whereas it remains stable in the redshift independent scenarios (Fig. 5 and Fig. 5) of the coalescence rate, as expected. For the redshift-dependent case, the number of sources at low redshift in high-frequency bins becomes so small that the expected number of sources in that frequency bin at that redshift is below 1. However, the Monte Carlo sampling only returns the discrete integer source. Therefore, the analytical curve lies below the simulated results at high frequencies. The effect of this is also seen in the $C_{\ell}(f)$ for two cases: when $\delta_{f}=0$ and $\mu=0$, and when $\delta_{f}=1$, $\mu=1$ and $z_{0}=1$, where there is a sudden jump at $f=2.4\times 10^{-8}$. Just like in the $\nu=0.5$ curve in Fig. 7, the expectation value of the number of sources in the low redshift bins, emitting at high frequencies bins, is less than one, but the sampled number of sources can be mostly either zero or one. This sporadic nature of the sampling at high-frequency bins results in a sharp increase in the anisotropy at high frequencies.
@@
The fitted curves for $C_{\ell}(f)$ show a significantly larger power law index ($\lambda$) for the redshift-dependent case compared to the redshift-independent case for the same $\delta_{f}$ value. This difference arises because the number of sources in the redshift-dependent case decreases as the frequency increases. Similarly, the power law index is larger for $\delta_{f}=0$ than that of $\delta_{f}=1$ for the redshift independent case because there are fewer sources at high frequencies for the former case than for the latter.
@@
The redshift evolution of the coalescence rate can help us understand the merger history of the galaxies. A consistent coalescence rate across redshifts suggests a stable merger rate of galaxies over time. Conversely, if relatively more binaries are found at lower frequencies at some redshifts, it could indicate a higher merger rate during those periods. In the case where the coalescence rate at low frequency (compared to high frequency) increases with redshift, signifies a larger galaxy merger rate at high redshifts.
@@
Variation of $C_{\ell}$ with spatial frequencies ($\ell$) and spectral frequencies($f$): In Fig. 10 we depict the spectral and spatial structure of the anisotropies in the SGWB for a few different models. The figure shows the mode of $C_{\ell}(f)$ over 1000 realizations as a function of the multipole moment ($\ell$) and frequency ($f$) in color. The results for different variations of the model parameters for each of these cases at a fixed spatial frequency $\ell=15$ are discussed previously. Here we mainly focus on the additional gain by measuring the signal in $\ell$ as well. The $\ell=0$ mode signifies the isotropic component ($\Omega_{\rm gw}(f)$) of the SGWB signal. The presence of non-zero higher-order moments implies additional information in the $\ell-f$ space beyond just the $\ell=0$ (isotropic) moment. The non-zero higher-order moments indicate the presence of anisotropy in the SGWB as a function of GW frequency. As evident from the color matrix plots in Fig. 10, the $\ell-f$ structure remains nearly invariant with respect to $\ell$ at all frequencies ($f$). This is expected as at these large angular scales, the properties of the host galaxies and the correlation with the n-Hz GW signal will be similar for a statistically isotropic and homogeneous Universe.
@@
In Fig. 10– Fig. 10, we display the $\ell-f$ plot of the $C_{\ell}(f)$ for four different cases. Consistent with the behavior at $\ell=15$, the $C_{\ell}(f)$ curves exhibit an increase as the frequency increases from $f=2\times 10^{-9}$ Hz to $f=3\times 10^{-8}$ Hz in nearly a steady way at all values of $\ell$. The steady increase at all $\ell$ with frequency $f$ makes the theoretical signal strongly correlated and predictable. As the Universe is expected to be statistically isotropic and homogeneous at large scales, the SGWB fluctuations at large scale are nearly independent of spatial scale $\ell$ values and vary strongly with the spectral frequency $f$ of GWs, which captures the evolution of the SMBHBs. If one can probe angular scales corresponding to galaxies (higher $\ell$s), the spatial behavior is expected to show some variation. As the current PTA analysis can only explore large angular, we have not explored the signal at small angular scales in this paper. The small angular scale signal will be explored in a future work.
@@
The presence of anisotropic power spectrum with more power at high $f$ for every value of spatial frequency $\ell$ makes it possible to measure the amplitude $B$ and the power-law index $\lambda$ (of a two-parameter model to capture the anisotropic power spectrum) of $C_{\ell}(f)$ by combining different values of $\ell$, which gives a $\sqrt{f_{\rm sky}(2\ell+1)}$ independent information per $\ell$ for a survey with sky-fraction $f_{\rm sky}$. The theoretical analysis presented in this paper demonstrates how the anisotropic signal of the SGWB in the $\ell-f$ space can capture the imprints of SMBHBs evolution in the Universe. We have shown the variation of $C_{\ell}(f)$ for different scenarios of astrophysical models and have theoretical predictions on the scaling of the anisotropic signal with $f$. In future work, we will explore the measurability of this signal and how it can be constrained from the SGWB observations.
@@
A summary of the key signatures of SMBHB evolution on the SGWB spatial-spectral frequency is as follows.
@@
  • Even in the absence of any redshift evolution of the SMBH mass and galaxy stellar mass relation, the anisotropy in SGWB captured in terms of the spatial-spectral power spectrum $C_{\ell}(f)$ can be large if SMBH masses and galaxy stellar mass relation is more steep implying that heavier SMBHs can be efficiently hosted in galaxies of lower stellar mass. The amplitude and shape of SGWB power spectrum $\Omega_{\rm gw}(f)$ can be comparable for both steep and shallow SMBH masses and galaxy stellar mass relation.
@@
  • If heavier SMBHs form efficiently at high redshift, then the amplitude of the anisotropy can increase in the high frequencies of n-Hz signal than the low frequency for a comparable SGWB power spectrum $\Omega_{\rm gw}(f)$. The spectral shape of $\Omega_{\rm gw}(f)$ and spatial-spectral anisotropy power spectra $C_{\ell}(f)$ bring complementary information on the redshift evolution of the SMBH and galaxy stellar mass relation.
@@
  • The coalescing rate of SMBHBs across cosmic time shows prominent features in both power spectrum $\Omega_{\rm gw}(f)$ and spatial-spectral anisotropy power spectrum $C_{\ell}(f)$. If the coalescence rate is less efficient in high frequency than in low frequency, and/or is less efficient at low redshift, then the amplitude of $C_{\ell}(f)$ increases. By measuring $\Omega_{\rm gw}(f)$ and $C_{\ell}(f)$ the frequency and redshift evolution of the coalescing rate can be inferred using the n-Hz signal.
Measuring the temporal fluctuation of the SGWB and estimate R(z) and mass distribution using temporal fluctuation of the SGWB .
@@ 7 Conclusion
However, to truly comprehend the SMBH population and its evolutionary trends, it is important to explore additional observational avenues and develop comprehensive theoretical frameworks that can account for the diverse range of processes involved in their formation and growth. Our study has highlighted the importance of studying the isotropic as well as the anisotropic SGWB density in the study of the evolution of SMBH which can shed light on the formation scenarios of SMBHBs. The study of SGWB anisotropy holds great promise as a valuable tool for advancing our understanding of some of the open questions regarding the formation and evolution of SMBHs and SMBHBs. The evolutionary history of the SMBHs is expected to have an imprint on the correlation between the SMBH mass and the properties of the host galaxies. To capture the physics from multi-scale dynamics of the SMBHB population, we adopted a multiscale adaptive technique to cross-match the physics involved at different scales from cosmological to the sub-pc scale where the coalescence of the binary occurs, to calculate the large-scale spatial and spectral anisotropy of the SGWB in n-Hz range. By utilizing large-scale simulations such as MICECAT and smaller-scale, high-resolution simulations such as ROMULUS25, along with an analytical model for the physics of the sub-pc SMBH environment, we demonstrate the impact of SMBHB evolution on the SGWB signal.
Our results have illustrated how the anisotropy of the SGWB from the population of the SMBHBs is affected by the SMBH mass-stellar mass relation. Similarly, we have also demonstrated that the coalescence rate and its evolution affect the overall shape of the spectral dependence of the angular correlation $C_{\ell}(f)$ of the SGWB. The coalescence rate of the SMBHBs at different separations of binaries can be used as a probe to the merger history of the galaxy. Furthermore, our investigation highlights the significant impact of environmental effects on the anisotropy of the SGWB. The presence of environmental effects, particularly at large separations of the binary, can alter the residence time of a binary at low frequencies. This effect leads to an increase in the anisotropy.
@@
Although with the present n-Hz data, the sky localization is poor, the future inclusion of the Square Kilometre Array (SKA) will significantly improve the sky localization error, as a large number of pulsars can be timed with SKA (Lazio, 2013; Stappers et al., 2018). SKA is projected to find several hundred pulsars, which will enhance the PTA’s ability to detect and localize the GW signal. With the addition of hundreds of pulsars, SKA is expected to enhance the localization of the GW signal to a few tens of square degrees (Sesana & Vecchio, 2010a, b).
The approach proposed in the work will help us put some valuable constraints on the population and evolution of the SMBHs. Using the multiscale adaptive technique employed in this work, we can efficiently calculate the spatial-spectral anisotropy of the SGWB signal for a wide range of astrophysical scenarios of SMBH formation and evolution, which can be tested with n-Hz GWs. In the future, incorporating this framework into a Bayesian data analysis setup will enable the measurement of parameters (such as $\rho,\eta,\nu,\delta_{f},\kappa,\mu,z_{0}$) by using the spatial-spectral anisotropic signal, $C_{\ell}(f)$. Furthermore, cross-correlation of the SGWB signal with the tracers of the large-scale structures (such as galaxies, quasars, AGN) will bring additional insights into the formation of the SMBHs in the Universe. A work is currently under preparation to explore these aspects. Though this study strictly limits the power spectrum of the SGWB, signatures beyond the power spectrum, such as non-Gaussian signatures are also expected to be important due to the sporadic spatial distribution of the SMBHBs in the Universe. These will also bring more constraining power in measuring the SMBHB’s evolution parameters.
@@ Acknowledgments
Data Availability
The data underlying this article will be shared at the request to the corresponding author.
@@ Appendix A Fitting power law to the 15-year data release of NANOGrav
We fit the $\Omega_{\rm gw}(f)$ of the NANOGrav 15-yr data set to a power law using PTArcade code (Mitridate et al., 2023).
$\Omega_{\rm gw}(f)=A\times(f/f_{yr})^{\alpha},$ (18)
where A is the amplitude of the $\Omega_{\rm gw}(f)$ at $f=f_{yr}$, $f_{yr}$ = 1/yr, and $\alpha$ is the power law index. Fig. 11 shows the corner plot of A and $\alpha$ fitted to the NANOGrav 15-yr data set. The median values along with the uncertainties on parameters based on the 16th, and 84th percentiles are $\rm Log_{10}$= $-7.14_{-0.29}^{+0.28}$ and $\alpha$= $1.76_{-0.34}^{+0.36}$.
In the cases: SMBH mass-Stellar mass and SMBH mass-Stellar mass evolution, we normalize the number of sources contributing to the signal in each frequency bin such that the mode of $\Omega_{\rm gw}(f)$ (over $10^{3}$ realizations) agrees with the power law curve obtained by taking the median values of the parameters A and $\alpha$.
Figure 11: Corner plot showing the posterior distributions of $\rm Log_{10}(A)$ and $\alpha$ for the power-law model fitted to $\Omega_{\rm gw}(f)$ estimates from NANOGrav 15-yr data set.